OpenStudy (anonymous):

What is the interquartile range of the following data set? 78, 90, 456, 676, 111, 381, 21

OpenStudy (darkbluechocobo):

First of all, you need to arrange the numbers in ascending order: 21, 78, 90, 111, 381, 456, 676 From here, you can see that median = 111 (middle number). After removing 111 from the set, you have two distinct sets: 1) 21, 78, 90 - from here we find Q1 = 78 (middle number) 2) 381, 456, 676 - from here we find Q3 = 456 (middle number) IQR = Q3 - Q1 = 456 - 78 = 378